let the random variable x represent the number of boys in the family construct the probability distribution for the family of two children





1
Expert's answer
2022-03-03T17:12:00-0500

"P(X=0)=(\frac{1}{2})^2=\frac{1}{4}."

"P(X=1)=(\frac{1}{2})^2*2=\frac{1}{2}."

"P(X=2)=(\frac{1}{2})^2=\frac{1}{4}."